### v1

# Cooperative Bistatic ISAC Systems for Low-Altitude Economy

Abstract
The burgeoning low-altitude economy (LAE) necessitates integrated sensing and communication (ISAC) systems capable of high-accuracy multi-target localization and velocity estimation under hardware and coverage constraints inherent in conventional ISAC architectures. This paper addresses these challenges by proposing a cooperative bistatic ISAC framework within MIMO-OFDM cellular networks, enabling robust sensing services for LAE applications through standardized 5G New Radio (NR) infrastructure. We first develop a low-complexity parameter extraction algorithm employing CANDECOMP/PARAFAC (CP) tensor decomposition, which exploits the inherent Vandermonde structure in delay-related factor matrices to efficiently recover bistatic ranges, Doppler velocities, and angles-of-arrival (AoA) from multi-dimensional received signal tensors. To resolve data association ambiguity across distributed transmitter-receiver pairs and mitigate erroneous estimates, we further design a robust fusion scheme based on the minimum spanning tree (MST) method, enabling joint 3D position and velocity reconstruction. Comprehensive simulation results validate the framework’s superiority in computational efficiency and sensing performance for low-altitude scenarios.
Index Terms:
Cooperative integrated sensing and communication (ISAC), orthogonal frequency division multiplexing (OFDM), low-altitude economy (LAE), bistatic sensing.
I Introduction
The rapid evolution toward sixth-generation (6G) and NextG wireless networks has ushered in a transformative vision where communication systems transcend traditional data transmission to embrace advanced sensing capabilities [1]. Integrated sensing and communication (ISAC) is central to realizing this vision, unifying sensing and communication within shared infrastructure. This convergence promises significant gains for optimized spectral efficiency, reduced hardware costs, and novel application domains [2]. By exploiting the inherent synergy between sensing and communication, ISAC not only enhances network adaptability but also paves the way for intelligent environments where real-time situational awareness coexists with ultra-reliable connectivity. ISAC is particularly transformative for the low-altitude economy (LAE), where its capabilities align directly with critical operational needs [3]. LAE encompasses applications spanning logistics, agriculture, and emergency services, which rely on unmanned aerial vehicles (UAVs) operating in complex airspace, demanding robust detection, localization, and tracking capabilities alongside seamless connectivity. Addressing these requirements necessitates ISAC protocols tailored to LAE’s unique challenges, including high-accuracy multi-target sensing and compatibility with the existing cellular infrastructure.
Similar to radar systems, ISAC implementations typically adopt two basic operational modes: monostatic (a.k.a. active sensing) [4, 5] and bistatic (a.k.a. passive sensing) [6]. In monostatic systems, a single base station (BS) simultaneously transmits and receives sensing signals, enabling independent operation. However, this approach requires full-duplex transceivers with self-interference mitigation capability, which is not supported by the existing commercial BSs, necessitating costly hardware upgrades. Conversely, bistatic configurations separate the transmitter and receiver, enabling sensing through reflected signals. This sensing pattern broadens coverage and avoid costly full-duplex hardware, facilitating integration with the existing systems [7, 8], while introducing synchronization requirements between BSs. In addition to these challenges, conventional configurations-whether monostatic or bistatic-also have inherent limitations: sparse spatial coverage due to limited transmission power and environmental obstacles, speed estimation restricted to the radial component, and susceptibility to measurement outliers.
Cooperative ISAC architectures, inspired by multistatic radar principles, are expected to overcome these limitations [9, 10]. By leveraging distributed BSs to collaboratively transmit and receive sensing signals, cooperative ISAC systems amplify coverage, enhance measurement diversity, and improve robustness against environmental dynamics. In cooperative ISAC networks, multiple BSs collaboratively transmit and receive signals, generating rich spatial-temporal measurements. Sensing parameters such as time delays, Doppler shifts, and angles are extracted locally at BSs or centrally at a central processing unit (CPU). These parameters are then fused at the CPU for high-accuracy target localization and true velocity reconstruction. In our previous work [11], a cell-free-based cooperative bistatic ISAC framework was proposed. It revealed that such cooperation can be achieved through protocol-level innovations, such as staggered uplink/downlink scheduling across BSs, which repurposed existing orthogonal frequency-division multiplexing (OFDM) communication waveforms and infrastructure for cooperative sensing without hardware updates. This alignment with the standardized 5G New Radio (NR) frame structure demonstrated how cooperative ISAC combines theoretical advancements with practical deployment capabilities, which is a key consideration for 6G standardization.
Prior research has established foundational frameworks for cooperative ISAC-enabled sensing schemes. For example, [11] and [12] addressed single-antenna BS configurations, employing bistatic and monostatic sensing modes through two-dimensional (2D) fast Fourier transform (2D-FFT) and compressed sensing algorithms for target localization. Subsequent studies extended these concepts to multi-antenna architectures. For the scenario with multi-antenna BSs, Lu et al. [13] investigated a multi-BS cooperative system using multiple signal classification (MUSIC) algorithms for parameter estimation, complemented by a lattice point search method for data fusion. This progression continued with Wei et al. [14], who developed a symbol-level data fusion scheme integrating phase features in the demodulation symbols with distance and radial velocity estimations in cooperative monostatic ISAC systems, later adapted by the same research group [15] for bistatic configurations with enhanced position and velocity estimation capabilities based on angle-of-arrival (AoA) and angle-of-departure (AoD) extractions. In addition, feasible topology and frame structure designs were proposed by Han et al. [16], which achieved cooperative bistatic sensing in the cellular networks at a low cost. A hybrid monostatic-bistatic architecture was proposed by Jiang et al. [17], where dual BSs jointly process echo signals from both BSs for improved target localization. Despite these advances, most studies focused on simplified scenarios, such as single-antenna configurations [12, 11], single-target sensing [13, 14, 15, 16], and 2D spatial estimation [12, 11, 14, 15, 17, 16], which failed to reflect the complexity of practical deployments like LAE.. Emerging solutions attempted to address these constraints through advanced mathematical tools. Tang et al. [18] introduced tensor decomposition techniques to cooperative monostatic systems, enabling low-complexity parameter estimation of range, angle, and velocity with inherent multi-dimensional parameter pairing. This approach has already been validated in channel estimation [19, 20], showing promise for the ISAC-enabled sensing scheme designs [21].
Given the extensive coverage of mobile networks, developing sensing schemes compatible with existing infrastructure is crucial for practical ISAC deployment and application proliferation. However, two key limitations hinder practical deployment: First, as mentioned above, despite the prevalence of massive MIMO BSs, existing studies predominantly addressed simplified single-antenna [12, 11] or single-target scenarios [13, 14, 15, 16], failing to leverage modern network capabilities. Second, bistatic architectures [11, 15, 16] have inherent advantages over monostatic approaches [12, 14, 13, 18] in hardware compatibility by reusing the existing communication equipment, while also generating richer sensing data. For instance, with 5 cooperative BSs configured as 2 transmitters and 3 receivers, bistatic sensing schemes can generate $2\times 3$ measurement links, surpassing the 5 self-transceiving links in monostatic schemes. Meanwhile, the reduction in the number of transmitters also decreases spectrum consumption and inter-BS interference. Nevertheless, fusing abundant bistatic measurements to enhance sensing accuracy remains challenging. While [11] proposed a solution for single-antenna scenarios, its high complexity limits its scalability in environments with a large number of BSs or targets. Although the data fusion framework in [18] serves as a viable paradigm for monostatic systems, it cannot be directly applied to bistatic architectures due to the inherent challenges in bistatic sensing, such as the elimination of symbol timing offset (STO) and carrier frequency offset (CFO), as well as issues related to three-dimensional (3D) position estimation.
This paper addresses the aforementioned gaps by developing a cooperative bistatic ISAC framework for LAE scenarios. Within a MIMO-OFDM cellular network, we introduce a bistatic sensing architecture compatible with standard 5G NR frame structures. To achieve high-accuracy position and velocity estimation, efficient signal processing methods are designed for sensing parameter extraction and data fusion. The main contributions of this work are summarized as follows:
  1. 1.
    We propose a general cooperative bistatic ISAC framework based on MIMO-OFDM cellular networks that enable multi-BS multi-target cooperative sensing with minimum hardware modifications, maintaining compatibility with 5G NR frame structures.
  2. 2.
    We develop a low-complexity parameter estimation algorithm using ESPRIT-inspired CANDECOMP/PARAFAC (CP) tensor decomposition, which efficiently extracts range, velocity, and angle parameters of the targets by leveraging the Vandermonde structure of delay-related factor matrices.
  3. 3.
    We design a robust data fusion scheme that employs a minimum spanning tree (MST)-based association method to eliminate erroneous estimates and resolve data association ambiguity among transmitter-receiver pairs, thereby enabling joint 3D position and velocity estimation.
  4. 4.
    Comprehensive simulation results validate the superiority of the proposed framework over the baseline schemes in terms of computational complexity and sensing performance in LAE scenarios.
The remainder of this paper is organized as follows. Section II introduces the system model and bistatic channel formulation. Section III details the tensor-based parameter extraction algorithm with uniqueness analysis. Section IV presents the MST-based data fusion framework. Section V evaluates the performance of the proposed schemes through simulations, and Section VI concludes the work.
Notations: Mathematical entities are typographically distinguished as follows: scalars use lowercase letters ($y$), vectors bold lowercase ($\mathbf{y}$), matrices bold uppercase ($\mathbf{Y}$), sets calligraphic uppercase ($\mathcal{Y}$), and tensors bold calligraphic uppercase ($\bm{\mathcal{Y}}$). The real and complex number fields are denoted by $\mathbb{R}$ and $\mathbb{C}$, respectively. Fundamental operators are defined as: $\circ$ for outer product, $\otimes$ for Kronecker product, * for Hadamard product, and $\odot$ for Khatri-Rao product. The transpose, conjugate transpose, inverse, and Moore-Penrose pseudo-inverse of a matrix $\mathbf{Y}$ are denoted by $\mathbf{Y}^{\mathrm{T}}$, $\mathbf{Y}^{\mathrm{H}}$, $\mathbf{Y}^{-1}$, and $\mathbf{Y}^{\dagger}$, respectively. The $\ell_{1}$ norm and $\ell_{2}$ norm of a vector $\mathbf{y}$ are expressed as $\left\|\mathbf{y}\right\|_{1}$ and $\left\|\mathbf{y}\right\|_{2}$, respectively. The operator $\mathrm{diag}\left(\mathbf{y}\right)$ generates a diagonal matrix from vector $\mathbf{y}$, while $\text{unvec}_{M\times N}\left(\cdot\right)$ reshapes an $MN\times 1$ vector $\mathbf{y}$ into an $M\times N$ matrix $\mathbf{Y}$.
II System Model
II-A Signal Transmission Model
As illustrated in Fig. 1, we consider a MIMO-OFDM network-based cooperative ISAC system capable of detecting, locating, and measuring velocities of low-altitude targets, thereby supporting the low-altitude economy applications. This system employs a bistatic sensing architecture where $N_{\mathrm{t}}$ transmit BSs (tBSs) simultaneously emit sensing signals, while $N_{\mathrm{r}}$ receive BSs (rBSs) collaboratively estimate the parameters of $K$ targets through processing the reflected echoes.111 While ISAC systems can dynamically reconfigure tBS/rBS assignments based on specific sensing requirements, this work focuses exclusively on developing sensing algorithms, with BS selection strategies proposed in [22]. Operating in the time division duplexing (TDD) mode, all BSs are connected with a CPU through fronthaul links. The CPU sends timing information to the BSs via the fronthaul links for synchronization, performs data fusion of the rBS-derived sensing information to enhance the sensing performance, and adjusts the slot format [23] of each BS to facilitate the sensing signal exchange. A detailed example of the slot format configuration is presented in Fig. 2. Similar to the cooperative ISAC framework established in our previous work [11], to achieve bistatic sensing in mobile communication systems, the rBSs are configured with slot formats different from the tBSs when performing sensing tasks. During the $N$-symbol sensing task, each tBS transmits pre-determined sensing reference signals on non-overlapping bandwidth parts (BWPs) to eliminate the inter-BS interference, enabling unambiguous signal separation and identification at all rBSs. The estimated parameters from these signals are then uploaded to the CPU through the fronthaul links.
Figure 1: An illustration of the MIMO-OFDM cellular network-based bistatic ISAC architecture.
Figure 2: An example of feasible slot format configuration, where a duration of $N=8$ symbols can be used for sensing signal.
Within the common 5G telecommunication infrastructure, every BS implements a MIMO-OFDM air interface employing uniform planar arrays (UPAs) for 3D beamforming. Specifically, the UPAs are installed with an elevation angle of 0° to facilitate both the sensing of low-altitude targets and the communication service for terrestrial users. To achieve cost-efficient massive antenna deployment, we consider an architecture that all BSs adopt the full-connected hybrid beamforming structure, which reduce the number of radio frequency (RF) chains. We assume that each BS comprises $R$ RF chains driving $L=N_{\mathrm{h}}\times N_{\mathrm{v}}>R$ antenna elements, where $N_{\mathrm{v}}$ and $N_{\mathrm{h}}$ denote the numbers of antenna elements in the vertical and horizontal directions, respectively.
In the considered sensing framework, all rBSs independently execute identical processing algorithms on signals transmitted from individual tBSs. In the following, we consider the signal received at rBS $n_{\mathrm{r}}$ as an illustrative case for presenting the transmission model and the proposed parameter extraction scheme. For notational compactness, we omit the subscripts $n_{\mathrm{t}}$ and $n_{\mathrm{r}}$ throughout the derivations in the following. Let $s_{l,m,n}$ denote the modulated frequency-domain sensing symbol allocated in the $m$th subcarrier and the $n$th OFDM symbol at the $l$th transmit antenna, which satisfies $\left|s_{l,m,n}\right|^{2}=1$. Then, the signal transmitted by tBS $n_{\mathrm{t}}$ can be represented as
$s_{l}\left(t\right)=\sum_{m=1}^{M}{\sum_{n=1}^{N}{s_{l,m,n}}e^{j2\pi m\Delta ft}}\cdot r\left(t-nT_{\mathrm{s}}\right),\,l=1,...,L,$ (1)
where $M$ is the number of subcarriers, $T_{\mathrm{s}}$ denotes the OFDM symbol period (including the cyclic prefix), $r(t)$ denotes the pulse-shaping filter function, and $\Delta f$ denotes the subcarrier spacing (SCS). Following matched filtering, cyclic prefix removing, and inverse fast Fourier transform (IFFT), the frequency-domain signal vector received at rBS $n_{\mathrm{r}}$ during the $n$th sensing symbol is expressed as
$\mathbf{y}_{m,n}=\mathbf{Q}^{\mathrm{H}}\mathbf{H}_{m,n}\mathbf{Fs}_{m,n}+\mathbf{Q}^{\mathrm{H}}\mathbf{n}_{m,n}\in\mathbb{C}^{R\times 1},$ (2)
where $\mathbf{H}_{m,n}\in\mathbb{C}^{L\times L}$ denotes the discrete frequency-domain channel (its formulation will be derived in the subsequent subsection), the hybrid precoding matrix $\mathbf{F}\triangleq\mathbf{F}_{\mathrm{a}}\mathbf{F}_{\mathrm{d}}$ combines an analog phase-shifting network $\mathbf{F}_{\mathrm{a}}\in\mathbb{C}^{L\times R}$ with a digital precoder $\mathbf{F}_{\mathrm{d}}\in\mathbb{C}^{R\times R}$, while the hybrid combining matrix $\mathbf{Q}=\mathbf{Q}_{\mathrm{a}}\mathbf{Q}_{\mathrm{d}}$ integrates an analog combining matrix $\mathbf{Q}_{\mathrm{a}}\in\mathbb{C}^{L\times R}$ and a digital combining matrix $\mathbf{Q}_{\mathrm{d}}\in\mathbb{C}^{R\times R}$. The transmitted symbol vector $\mathbf{s}_{m,n}=[s_{1,m,n},\cdots,s_{l,m,n},\cdots,s_{L,m,n}]^{\mathrm{T}}\in\mathbb{C}^{L\times 1}$ contains the sensing symbols from tBS $n_{\mathrm{t}}$, and $\mathbf{n}_{m,n}\in\mathbb{C}^{L\times 1}$ represents the additive white Gaussian noise vector. Given the known symbol vector $\mathbf{s}_{m,n}$, the rBSs can eliminate the impact of it by multiplying the received signal vector by $\mathbf{s}_{m,n}^{*}$, i.e.,
$\hat{\mathbf{y}}_{m,n}=\mathbf{s}_{m,n}^{*}*\mathbf{y}_{m,n}=\mathbf{Q}^{\mathrm{H}}\mathbf{H}_{m,n}\mathbf{F1}+\hat{\mathbf{n}}_{m,n},$ (3)
where $\mathbf{1}$ is an all-one column vector and $\hat{\mathbf{n}}_{m,n}\in\mathbb{C}^{L\times 1}$ is the equivalent noise vector.
II-B Sensing Channel Model
In low-altitude sensing scenarios, consistent with common assumptions in cooperative ISAC systems [11, 18], we assume a direct line-of-sight (LoS) path between any tBS-rBS pair. This leads to $K+1$ resolvable propagation paths at rBS $n_{\mathrm{r}}$, comprising $K$ target-scattered paths and one dominant LoS component, where the length of target-scattered paths are also known as the bistatic ranges. The temporal characteristics of these paths are quantified through path-specific delays $\tau_{k}=d_{k}/c_{0}$ for $k=1,\dots,K+1$, where $d_{k}$ represents the length of the $k$th path, and $c_{0}$ is the speed of light. In addition, the Doppler shift of the LoS path ($k=k_{\mathrm{LoS}}$) is known to satisfy $f_{k_{\mathrm{LoS}}}^{\mathrm{d}}=0$, while each scattered path ($k\neq k_{\mathrm{LoS}}$) exhibits shift $f_{k}^{\mathrm{d}}=v_{k}/\lambda$ proportional to the total radial velocity $v_{k}$ (called bistatic Doppler velocity) of a target relative to tBS $n_{\mathrm{t}}$ and rBS $n_{\mathrm{r}}$, with $\lambda$ denoting the carrier wavelength. Then, the time-delay domain channel matrix between tBS $n_{\mathrm{t}}$ and rBS $n_{\mathrm{r}}$ is formulated as
$\mathbf{H}(t,\tau)=\sum_{k=1}^{K+1}{\left[\begin{array}[]{l}\alpha_{k}\mathbf{a}_{\mathrm{r}}\left(\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\right)\mathbf{a}_{\mathrm{t}}^{\mathrm{H}}\left(\theta_{k}^{\mathrm{t}},\phi_{k}^{\mathrm{t}}\right)\\ \times\delta\left(\tau-\tau_{k}-{\delta}^{\mathrm{t}}\right)e^{j2\pi(f_{k}^{\mathrm{d}}+{\delta}^{\mathrm{f}})t}\\ \end{array}\right],}$ (6)
where $\alpha_{k}$ represents the composite channel gain of the $k$th path, while ${\delta}^{\mathrm{t}}$ and ${\delta}^{\mathrm{f}}$ are the STO and CFO after synchronization between tBS $n_{\mathrm{t}}$ and rBS $n_{\mathrm{r}}$. The spatial characteristics are determined by azimuth/elevation angles of arrival (AoA) $\left\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\right\}$ and departure (AoD) $\left\{\theta_{k}^{\mathrm{t}},\phi_{k}^{\mathrm{t}}\right\}$. To simplify the notation, we further define
$\left\{\begin{array}[]{l}\varPhi_{k}^{\mathrm{r}}\triangleq\cos\theta_{k}^{\mathrm{r}},\,\varTheta_{k}^{\mathrm{r}}\triangleq\sin\theta_{k}^{\mathrm{r}}\cos\phi_{k}^{\mathrm{r}},\\ \varPhi_{k}^{\mathrm{t}}\triangleq\cos\theta_{k}^{\mathrm{t}},\,\varTheta_{k}^{\mathrm{t}}\triangleq\sin\theta_{k}^{\mathrm{t}}\cos\phi_{k}^{\mathrm{t}}.\\ \end{array}\right.$ (7)
Then, the steering vectors $\mathbf{a}_{\mathrm{r}}\left(\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\right)$ and $\mathbf{a}_{\mathrm{t}}\left(\theta_{k}^{\mathrm{t}},\phi_{k}^{\mathrm{t}}\right)$ in (6) can be written as
$\mathbf{a}_{\mathrm{r}}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}}\right)=\mathbf{a}_{\mathrm{r}}^{\mathrm{v}}\left(\varPhi_{k}^{\mathrm{r}}\right)\otimes\mathbf{a}_{\mathrm{r}}^{\mathrm{h}}\left(\varTheta_{k}^{\mathrm{r}}\right)\in\mathbb{C}^{L\times 1},$ (8)
$\mathbf{a}_{\mathrm{t}}\left(\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)=\mathbf{a}_{\mathrm{t}}^{\mathrm{v}}\left(\varPhi_{k}^{\mathrm{t}}\right)\otimes\mathrm{}\mathbf{a}_{\mathrm{t}}^{\mathrm{h}}\left(\varTheta_{k}^{\mathrm{t}}\right)\in\mathbb{C}^{L\times 1},$ (9)
where the corresponding UPA steering vectors are constructed based on planar wavefront geometry
$\mathbf{a}_{\mathrm{r}}^{\mathrm{v}}\left(\varPhi_{k}^{\mathrm{r}}\right)=[1,\cdots,e^{j2\pi(N_{\mathrm{v}}-1)\frac{d}{\lambda}\varPhi_{k}^{\mathrm{r}}}]^{\mathrm{T}}\in\mathbb{C}^{N_{\mathrm{v}}\times 1},$ (10)
$\mathbf{a}_{\mathrm{r}}^{\mathrm{h}}\left(\varTheta_{k}^{\mathrm{r}}\right)=[1,\cdots,e^{j2\pi(N_{\mathrm{h}}-1)\frac{d}{\lambda}\varTheta_{k}^{\mathrm{r}}}]^{\mathrm{T}}\in\mathbb{C}^{N_{\mathrm{h}}\times 1},$ (11)
with analogous expressions for $\mathbf{a}_{\mathrm{t}}^{\mathrm{v}}\left(\varPhi_{k}^{\mathrm{t}}\right)$ and $\mathbf{a}_{\mathrm{t}}^{\mathrm{h}}\left(\varTheta_{k}^{\mathrm{t}}\right)$ through substitution of superscripts $\mathrm{r}\rightarrow\mathrm{t}$. By processing the Fourier transform of $\tau$, the frequency domain channel at the $m$th subcarrier can be derived as
$\mathbf{H}_{m}(t)=\sum_{k=1}^{K+1}{\left[\begin{array}[]{l}\alpha_{k}\mathbf{a}_{\mathrm{r}}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}}\right)\mathbf{a}_{\mathrm{t}}^{\mathrm{H}}\left(\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)\\ \times e^{-j2\pi m\Delta f\left(\tau_{k}+{\delta}^{\mathrm{t}}\right)}e^{j2\pi\left(f_{k}^{\mathrm{d}}+{\delta}^{\mathrm{f}}\right)t}\\ \end{array}\right].}$ (14)
Then, by sampling the received signal at the $n$th OFDM symbol, we can obtain the discrete frequency-domain channel for the $n$th symbol at the $m$th subcarrier as follows
$\mathbf{H}_{m,n}=\sum_{k=1}^{K+1}{\left[\begin{array}[]{l}\alpha_{k}\mathbf{a}_{\mathrm{r}}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}}\right)\mathbf{a}_{\mathrm{t}}^{\mathrm{H}}\left(\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)\\ \times e^{-j2\pi m\Delta f\left(\tau_{k}+{\delta}^{\mathrm{t}}\right)}e^{j2\pi\left(f_{k}^{\mathrm{d}}+{\delta}^{\mathrm{f}}\right)nT_{\mathrm{s}}}\\ \end{array}\right].}$ (17)
In the following, we first develop an efficient sensing parameter estimation algorithm to extract the bistatic ranges $\left\{d_{k}\right\}_{k=1}^{K+1}$, azimuth/elevation angles $\left\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\right\}_{k=1}^{K+1}$, and bistatic Doppler velocities $\left\{v_{k}\right\}_{k=1}^{K+1}$ related to the targets from the received signals $\hat{\mathbf{y}}_{m,n}$. Then, we propose a data fusion scheme that integrates these parameters to achieve accurate 3D localization and velocity estimation of the targets.
III Tensor-Based Parameter Estimation Algorithm
In this section, a tensor decomposition-based parameter estimation framework for bistatic ISAC systems is proposed. By reshaping received signals into a third-order tensor, we employ CP decomposition to extract the delay, angle, and Doppler factor matrices from the received signals $\hat{\mathbf{y}}_{m,n}$. Leveraging the inherent Vandermonde structure in one factor matrix, we develop an ESPRIT-inspired algorithm for efficient parameter estimation.
III-A Tensor Formulation of the Received Signal
By substituting (17) into (3), the received signal $\hat{\mathbf{y}}_{m,n}$ can be rewritten as
$\hat{\mathbf{y}}_{m,n}=\sum_{k=1}^{K+1}{\left[\begin{array}[]{l}\alpha_{k}\mathbf{Q}^{\mathrm{H}}\mathbf{a}_{\mathrm{r}}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}}\right)\mathbf{a}_{\mathrm{t}}\left(\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)^{H}\mathbf{F1}\\ \times e^{-j2\pi m\Delta f\left(\tau_{k}+{\delta}^{\mathrm{t}}\right)}e^{j2\pi\left(f_{k}^{\mathrm{d}}+{\delta}^{\mathrm{f}}\right)nT_{\mathrm{s}}}\\ \end{array}\right]}+\hat{\mathbf{n}}_{m,n}.$ (20)
By stacking the $N$ OFDM symbols and $M$ subcarriers, the received signal can be transformed to a three-order tensor $\bm{\mathcal{Y}}\in\mathbb{C}^{R\times N\times M}$ as follows
$\bm{\mathcal{Y}}=\sum_{k=1}^{K+1}{\alpha_{k}\mathbf{a}(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}},\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}})\circ\mathbf{b}(f_{k}^{\mathrm{d}},{\delta}^{\mathrm{f}})\circ\mathbf{c}(\tau_{k},{\delta}^{\mathrm{t}})}+\mathcal{N},$ (21)
where $\mathcal{N}$ is the noise tensor, $\mathbf{a}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}},\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)\in\mathbb{C}^{R\ \times 1}$, $\mathbf{b}\left(f_{k}^{\mathrm{d}},{\delta}^{\mathrm{f}}\right)\in\mathbb{C}^{N\times 1}$, and $\mathbf{c}\left(\tau_{k},{\delta}^{\mathrm{t}}\right)\in\mathbb{C}^{M\times 1}$ can be respectively derived as follows
$\!\mathbf{a}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}},\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)$ $=\mathbf{Q}^{\mathrm{H}}\mathbf{a}_{r}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}}\right)\mathbf{a}^{\mathrm{H}}_{t}\left(\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)\mathbf{F1},\!$ (22)
$\mathbf{b}(f_{k}^{\mathrm{d}},{\delta}^{\mathrm{f}})$ $=[1,\cdots,e^{j2\pi(N-1)T_{\mathrm{s}}(f_{k}^{\mathrm{d}}+{\delta}^{\mathrm{f}})}]^{\mathrm{T}},$ (23)
$\mathbf{c}\left(\tau_{k},{\delta}^{\mathrm{t}}\right)$ $=[1,\cdots,e^{-j2\pi(M-1)\Delta f(\tau_{k}+{\delta}^{\mathrm{t}})}]^{\mathrm{T}}.$ (24)
By defining factor matrices $\mathbf{A}\in\mathbb{C}^{R\times(K+1)}$, $\mathbf{B}\in\mathbb{C}^{N\times(K+1)}$, and $\mathbf{C}\in\mathbb{C}^{M\times(K+1)}$, whose $k$th column are respectively given by $\mathbf{a}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}},\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)$, $\alpha_{k}\mathbf{b}(f_{k}^{\mathrm{d}},{\delta}^{\mathrm{f}})$, and $\mathbf{c}\left(\tau_{k},{\delta}^{\mathrm{t}}\right)$, (21) can also be expressed as
$\bm{\mathcal{Y}}=\llbracket\mathbf{A},\mathbf{B},\mathbf{C}\rrbracket+\mathcal{N}.$ (25)
III-B Decomposition-Based Factor Matrices Recovery
From (22)-(24), it can be observed that the factor matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ respectively encode the angle, Doppler, and delay parameters of all propagation paths. To extract these sensing parameters, we first formulate the following CP decomposition problem to recover $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$:
$\min_{\mathbf{A},\mathbf{B},\mathbf{C}}\left\|\bm{\mathcal{Y}}-\sum_{k=1}^{K+1}{\alpha_{k}\mathbf{a}(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}},\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}})\!\circ\!\mathbf{b}(f_{k}^{\mathrm{d}},{\delta}^{\mathrm{f}})\!\circ\!\mathbf{c}(\tau_{k},{\delta}^{\mathrm{t}})}\right\|_{\mathrm{F}}^{2},$ (26)
where $\|\cdot\|_{\mathrm{F}}$ denotes the Frobenius norm. While conventional alternating least squares (ALS) methods [19, 24] can solve the non-convex Problem (26), they suffer from computational inefficiency, initialization-dependent convergence, and numerical instability. To address these limitations, in the following, we leverage the inherent Vandermonde structure of delay-related matrix $\mathbf{C}$ through an ESPRIT-inspired decomposition scheme.
Let $\mathbf{Y}_{(1)}$ represent the mode-1 unfolding of tensor $\bm{\mathcal{Y}}$, expressed as
$\mathbf{Y}_{(1)}^{\mathrm{T}}=\left(\mathbf{C}\odot\mathbf{B}\right)\mathbf{A}^{\mathrm{T}}\in\mathbb{C}^{MN\times R}.$ (27)
In addition, we select a pair of integer ${L_{1},L_{2}}$ subjected to $L_{1}+L_{2}=M+1$ and define the cyclic matrix as
$\mathbf{J}_{l}=\left[\mathbf{0}_{L_{1}\times(l-1)},\mathbf{I}_{L_{1}},\mathbf{0}_{L_{1}\times(L_{2}-l)}\right]\in\mathbb{C}^{L_{1}\times M}.$ (28)
Then, a smoothed measurement matrix can be constructed as
$\mathbf{Y}_{\mathrm{S}}$ $=\left[\left(\mathbf{J}_{1}\otimes\mathbf{I}_{N}\right)\mathbf{Y}_{\left(1\right)}^{\mathrm{T}},\dots,\left(\mathbf{J}_{L_{2}}\otimes\mathbf{I}_{N}\right)\mathbf{Y}_{\left(1\right)}^{\mathrm{T}}\right]$
=$\overset{\left(a\right)}{=}\left(\mathbf{C}^{L_{1}}\odot\mathbf{B}\right)\left(\mathbf{C}^{L_{2}}\odot\mathbf{A}\right)^{\mathrm{T}}\in\mathbb{C}^{L_{1}N\times L_{2}R},$ (29)
where $\mathbf{C}^{L_{1}}=[\mathbf{C}]_{1:L_{1},:}$ and $\mathbf{C}^{L_{2}}=[\mathbf{C}]_{1:L_{2},:}$ denote the first $L_{1}$ and $L_{2}$ rows of $\mathbf{C}$, respectively. The equality $(a)$ follows from Khatri-Rao product properties, i.e., $(\mathbf{A}\otimes\mathbf{B})(\mathbf{C}\odot\mathbf{D})=(\mathbf{A}\mathbf{C})\odot(\mathbf{B}\mathbf{D})$, and the shift-invariance induced by the Vandermonde structure of $\mathbf{C}$, i.e., $\mathbf{J}_{l_{2}}\mathbf{C}=\mathbf{J}_{1}\mathbf{C}\mathrm{diag}\left(\left[\mathbf{C}\right]_{l_{2},:}\right)$ [25]. By performing truncated singular value decomposition (SVD) on $\mathbf{Y}_{\mathrm{S}}$, we have
$\mathbf{Y}_{\mathrm{S}}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\mathrm{H}},$ (30)
where $\mathbf{U}\in\mathbb{C}^{L_{1}N\times(K+1)}$, $\mathbf{\Sigma}\in\mathbb{C}^{(K+1)\times(K+1)}$, and $\mathbf{V}\in\mathbb{C}^{L_{2}R\times(K+1)}$. Note that $\mathbf{U}$ span the column spaces of $\mathbf{Y}_{S}$. Therefore, a nonsingular matrix $\mathbf{M}\in\mathbb{C}^{(K+1)\times(K+1)}$ exists such that [26]
$\mathbf{C}^{L_{1}}\odot\mathbf{B}$ $=\mathbf{UM}\in\mathbb{C}^{L_{1}N\times(K+1)},$ (31)
$\mathbf{C}^{L_{2}}\odot\mathbf{A}$ $=\mathbf{V}^{*}\mathbf{\Sigma}\left(\mathbf{M}^{-1}\right)^{\mathrm{T}}\in\mathbb{C}^{L_{2}R\times(K+1)}.$ (32)
By leveraging the shift-invariant structure, we partition $\mathbf{U}$ into overlapped submatrices $\mathbf{U}_{1}={\left[\mathbf{U}\right]}_{1:(L_{1}-1)N,:}$ and $\mathbf{U}_{2}=\left[\mathbf{U}\right]_{N+1:L_{1}N,:}$, yielding
$\mathbf{U}_{1}\mathbf{M}$ $=\underline{\mathbf{C}}^{L_{1}}\odot\mathbf{B},$ (33)
$\mathbf{U}_{2}\mathbf{M}$ $=\overline{\mathbf{C}}^{L_{1}}\odot\mathbf{B},$ (34)
where $\underline{\mathbf{C}}^{L_{1}}$ and $\overline{\mathbf{C}}^{L_{1}}$ respectively denote $\mathbf{C}^{L_{1}}$ with the first and last rows removed. Based on the Vandermonde structure of $\mathbf{C}^{L_{1}}$, we derive that
$(\underline{\mathbf{C}}^{L_{1}}\odot\mathbf{B})\mathbf{Z}=\overline{\mathbf{C}}^{L_{1}}\odot\mathbf{B},$ (35)
where $\mathbf{Z}=\mathrm{diag}([e^{-j2\pi\Delta f\Delta\tau_{1}},\ldots,e^{-j2\pi\Delta f\Delta\tau_{K}}])$. Substituting (33) and (34) into (35), we have
$\mathbf{U}_{1}^{\dagger}\mathbf{U}_{2}=\mathbf{MZM}^{-1}.$ (36)
From (36), it can be noted that the estimates of $\mathbf{Z}$ and $\mathbf{M}$, denoted by $\hat{\mathbf{Z}}$ and $\hat{\mathbf{M}}$, can be obtained from the eigenvalue decomposition (EVD) of $\mathbf{U}_{1}^{\dagger}\mathbf{U}_{2}$. Then, a recovery of $\mathbf{C}$ can be generated by the normalized diagonal elements of $\hat{\mathbf{Z}}$. Specifically, each column of the recovered factor matrix ${\mathbf{C}}$ is given by
$\hat{\mathbf{c}}_{k}=\left[1,\hat{z}_{k},\ldots,\hat{z}_{k}^{M-1}\right]^{\mathrm{T}},\,k=1,\ldots,K+1,$ (37)
where $\hat{z}_{k}={[\hat{\mathbf{Z}}]_{k,k}}\big{/}{\left|[\hat{\mathbf{Z}}]_{k,k}\right|}$. Recall that the $k$th column of $\mathbf{B}$ is given by $\alpha_{k}\mathbf{b}(f_{k}^{\mathrm{d}},{\delta}^{\mathrm{f}})$. Let $\mathbf{c}_{k}^{L_{1}}$ and $\mathbf{m}_{k}$ denote the $k$th columns of $\mathbf{C}^{L_{1}}$ and $\mathbf{M}_{k}$, respectively. From the Kronecker product structure in (31), we have
$\mathbf{c}_{k}^{L_{1}}\otimes\left(\alpha_{k}\mathbf{b}(f_{k}^{\mathrm{d}},{\delta}^{\mathrm{f}})\right)=\mathbf{Um}_{k}.$ (38)
Hence, we can further recover each column of $\mathbf{B}$ as follows
$\hat{\mathbf{b}}_{k}\overset{\left(a\right)}{=}$ $\left(\frac{(\hat{\mathbf{c}}_{k}^{L_{1}})^{\mathrm{H}}}{\left\|\hat{\mathbf{c}}_{k}^{L_{1}}\right\|_{2}^{2}}\otimes\mathbf{I}_{N}\right)\left(\hat{\mathbf{c}}_{k}^{L_{1}}\otimes\hat{\mathbf{b}}_{k}\right)$
$=$ $\left(\frac{(\hat{\mathbf{c}}_{k}^{L_{1}})^{\mathrm{H}}}{\left\|\hat{\mathbf{c}}_{k}^{L_{1}}\right\|_{2}^{2}}\otimes\mathbf{I}_{N}\right)\mathbf{U}\hat{\mathbf{m}}_{k},\,k=1,\cdots,K+1,$ (39)
where $\hat{\mathbf{c}}_{k}^{L_{1}}$ and $\hat{\mathbf{m}}_{k}$ represent the $k$th columns of $\hat{\mathbf{C}}^{L_{1}}$ and $\hat{\mathbf{M}}_{k}$, respectively. The equality $\left(a\right)$ in (39) follows from the Kronecker product identities, i.e., $\left(\mathbf{A}\otimes\mathbf{B}\right)\left(\mathbf{C}\otimes\mathbf{D}\right)=\mathbf{AC}\otimes\mathbf{BD}$. Similarly, it can be derived from (31) that $\mathbf{c}_{k}^{L_{2}}\otimes\mathbf{a}_{k}=\mathbf{V^{*}}\mathbf{\Sigma}\mathbf{t}_{k}$, where $\mathbf{t}_{k}$ is the $k$th column of $\left(\mathbf{M}^{-1}\right)^{\mathrm{T}}$. Thus, each column of ${\mathbf{A}}$ can be recovered as
$\hat{\mathbf{a}}_{k}=\left(\frac{(\hat{\mathbf{c}}_{k}^{L_{2}})^{\mathrm{H}}}{\left\|\hat{\mathbf{c}}_{k}^{L_{2}}\right\|_{2}^{2}}\otimes\mathbf{I}_{R}\right)\mathbf{V}^{*}\mathbf{\Sigma}\hat{\mathbf{t}}_{k},\,k=1,\cdots,K+1,$ (40)
where $\hat{\mathbf{c}}_{k}^{L_{2}}$ and $\hat{\mathbf{t}}_{k}$ represent the $k$th columns of $\hat{\mathbf{C}}^{L_{2}}$ and $(\hat{\mathbf{M}}^{-1})^{\mathrm{T}}$, respectively.
Remark 1
Here, the factor matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ are recovered as $\hat{\mathbf{A}}=\left[\hat{\mathbf{a}}_{1},\dots,\hat{\mathbf{a}}_{K+1}\right]$, $\hat{\mathbf{B}}=\left[\hat{\mathbf{b}}_{1},\dots,\hat{\mathbf{b}}_{K+1}\right]$, and $\hat{\mathbf{C}}=\left[\hat{\mathbf{c}}_{1},\dots,\hat{\mathbf{c}}_{K+1}\right]$, respectively. The uniqueness of the CP decomposition will be discussed in Section III-D, which guarantees the following equalities [27]:
$\hat{\mathbf{A}}=\mathbf{A}\mathbf{\Pi}\mathbf{\Delta}_{1},\ \hat{\mathbf{B}}=\mathbf{B}\mathbf{\Pi}\mathbf{\Delta}_{2},\ \hat{\mathbf{C}}=\mathbf{C}\mathbf{\Pi}\mathbf{\Delta}_{3},$ (41)
where $\mathbf{\Pi}$ is a unique permutation matrix, while $\mathbf{\Delta}_{1}$, $\mathbf{\Delta}_{2}$, and $\mathbf{\Delta}_{3}$ are unique diagonal scaling matrices satisfying $\mathbf{\Delta}_{1}\mathbf{\Delta}_{2}\mathbf{\Delta}_{3}=\mathbf{I}_{K+1}$. The shared permutation matrix ensures automatic parameter pairing - the $k$th columns across all factor matrices inherently correspond to the same propagation path, which reduces the data association requirements.
III-C Parameter Estimation
In this subsection, we extract the bistatic ranges, azimuth/elevation angles, and bistatic Doppler velocities correspond to all targets from the recovered factor matrices $\hat{\mathbf{A}}$, $\hat{\mathbf{B}}$, and $\hat{\mathbf{C}}$.
First, the delay of the $k$th path can be estimated as
$\Delta\hat{\tau}_{k}\triangleq\hat{\tau}_{k}+\hat{\delta}^{\mathrm{t}}=\frac{\measuredangle\hat{z}_{k}}{-2\pi\Delta f},\,k=1,\cdots,K+1,$ (42)
where $\hat{\delta}^{\mathrm{t}}$ denotes the estimate of the STO ${\delta}^{\mathrm{t}}$, and $\measuredangle\hat{z}_{k}$ is the angle of $\hat{z}_{k}$. Without estimating the permutation matrix $\mathbf{\Pi}$, we can index the minimum value of $\{\Delta\hat{\tau}_{k}\}_{k=1}^{K+1}$ to find the LoS path, i.e.,
$k_{\mathrm{LoS}}=\arg\min_{\begin{subarray}{c}1\leqslant k\leqslant K+1\end{subarray}}\Delta\hat{\tau}_{k}.$ (43)
Since the locations of all BSs are apriori known, the residual STO can be estimated as
$\hat{\delta}^{t}=\Delta\hat{\tau}_{k_{\mathrm{LoS}}}-{\tau}_{k_{\mathrm{LoS}}},$ (44)
where $\tau_{k_{\mathrm{LoS}}}=d_{k_{\mathrm{LoS}}}/c_{0}$, and $d_{k_{\mathrm{LoS}}}$ is the distance between tBS $n_{\mathrm{t}}$ and rBS $n_{\mathrm{r}}$. Then, the delay of the target-scattered paths can be extracted as
$\hat{\tau}_{k}=\Delta\hat{\tau}_{k}-\hat{\delta}^{t},\,k\in\{1,\cdots,K+1\}\setminus k_{\mathrm{LoS}}.$ (45)
The corresponding bistatic ranges are obtained as
$\hat{d}_{k}=\hat{\tau}_{k}c_{0},\,k\in\{1,\cdots,K+1\}\setminus k_{\mathrm{LoS}}.$ (46)
Next, the estimation of the Doppler parameters $\{f^{\mathrm{d}}_{k},{\delta}^{\mathrm{f}}\}$ can be deduced as
$\Delta\hat{f}_{k}^{\mathrm{d}}\triangleq\hat{\delta}^{\mathrm{f}}+\hat{f}_{k}^{\mathrm{d}}=\mathrm{arg}\max_{\Delta f_{k}^{\mathrm{d}}}\frac{\left|\hat{\mathbf{b}}_{k}^{\mathrm{H}}\mathbf{b}\left(f_{k}^{\mathrm{d}},\delta^{\mathrm{f}}\right)\right|^{2}}{\left\|\hat{\mathbf{b}}_{k}\right\|_{2}^{2}\left\|\mathbf{b}\left(f_{k}^{\mathrm{d}},\delta^{\mathrm{f}}\right)\right\|_{2}^{2}},$
$k\in\{1,\cdots,K+1\}\setminus k_{\mathrm{LoS}}.$ (47)
The solution of Problem (47) can be obtained via the one-dimensional search. By utilizing $f_{k_{\mathrm{LoS}}}^{\mathrm{d}}=0$, we can separate $\hat{f}_{k}^{\mathrm{d}}$ and $\hat{\delta}^{\mathrm{f}}$ as follows
$\hat{\delta}^{\mathrm{f}}=\Delta\hat{f}_{k_{\mathrm{LoS}}}^{\mathrm{d}},$ (48)
$\hat{f}_{k}^{\mathrm{d}}=\Delta\hat{f}_{k}^{\mathrm{d}}-\hat{\delta}^{\mathrm{f}},\,k\in\{1,\cdots,K+1\}\setminus k_{\mathrm{LoS}}.$ (49)
Then, the bistatic Doppler velocities are estimated as
$\hat{v}_{k}=\hat{f}^{\mathrm{d}}_{k}\lambda,\,k\in\{1,\cdots,K+1\}\setminus k_{\mathrm{LoS}}.$ (50)
Finally, we extract the angle parameters from $\hat{\mathbf{A}}$. Note that it is challenging for the four-dimensional parameter extraction of $\left\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}},\theta_{k}^{\mathrm{t}},\phi_{k}^{\mathrm{t}}\right\}_{k=1}^{K+1}$. Therefore, we propose to only estimate the AoAs $\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\}_{k=1}^{K+1}$. The AoDs $\{\theta_{k}^{\mathrm{t}},\phi_{k}^{\mathrm{t}}\}_{k=1}^{K+1}$ can be calculated after target localization. From (22), we formulate the problem for jointly estimating ${\varTheta}_{k}^{\mathrm{r}}$ and ${\varPhi}_{k}^{\mathrm{r}}$ as
$\max_{\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}}}\frac{\left|\hat{\mathbf{a}}_{k}^{\mathrm{H}}\mathbf{Q}^{\mathrm{H}}\mathbf{a}_{\mathrm{r}}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}}\right)\right|^{2}}{\left\|\hat{\mathbf{a}}_{k}\right\|_{2}^{2}\left\|\mathbf{Q}^{\mathrm{H}}\mathbf{a}_{\mathrm{r}}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}}\right)\right\|_{2}^{2}},$ (51)
where the normalization operation eliminates the scaling ambiguity caused by $\mathbf{\Delta}_{1}$ and the impact of AoDs. In the objective function of Problem (51), the variables ${\varTheta}_{k}^{\mathrm{r}}$ and ${\varPhi}_{k}^{\mathrm{r}}$ are highly coupled. Although the problem can be solved via a two-dimensional search, the complexity is prohibitively high. In the following, an SVD-based decoupling algorithm is proposed for estimating $\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\}_{k=1}^{K+1}$ efficiently. First, we approximate the combining matrix $\mathbf{Q}$ as the Kronecker product of two matrices by solving the following problem [28]:
$\min_{\mathbf{Q}_{\mathrm{v}},\mathbf{Q}_{\mathrm{h}}}\parallel\mathbf{Q}-\mathbf{Q}_{\mathrm{v}}\otimes\mathbf{Q}_{\mathrm{h}}\parallel_{\mathrm{F}}^{2},$ (52)
where $\mathbf{Q}_{\mathrm{v}}\in\mathbb{C}^{N_{\mathrm{v}}\times J_{\mathrm{v}}}$, $\mathbf{Q}_{\mathrm{h}}\in\mathbb{C}^{N_{\mathrm{h}}\times J_{\mathrm{h}}}$, and $J_{\mathrm{v}}J_{\mathrm{h}}=R$. Problem (52) can be equivalently reformulated as
$\min_{\mathbf{Q}_{\mathrm{v}},\mathbf{Q}_{\mathrm{h}}}\left\|R\left(\mathbf{Q}\right)-\mathrm{vec}\left(\mathbf{Q}_{\mathrm{v}}\right)\mathrm{vec}\left(\mathbf{Q}_{\mathrm{h}}\right)^{\mathrm{T}}\right\|_{\mathrm{F}}^{2},$ (53)
where $R\left(\mathbf{Q}\right)$ is defined as follows:
$R\left(\mathbf{Q}\right)=\left[\begin{array}[]{c}\mathbf{Q}_{1}\\ \vdots\\ \mathbf{Q}_{j_{\mathrm{v}}}\\ \vdots\\ \mathbf{Q}_{J_{\mathrm{v}}}\\ \end{array}\right],\ \mathbf{Q}_{j_{\mathrm{v}}}=\left[\begin{array}[]{c}\mathrm{vec}\left(\left[\mathbf{Q}\right]_{1,j_{\mathrm{v}}}\mathbf{Q}_{\mathrm{h}}\right)^{\mathrm{T}}\\ \vdots\\ \mathrm{vec}\left(\left[\mathbf{Q}\right]_{N_{\mathrm{v}},j_{\mathrm{v}}}\mathbf{Q}_{\mathrm{h}}\right)^{\mathrm{T}}\\ \end{array}\right].$ (54)
The rank-one approximation problem (53) can be solved by performing SVD on $R\left(\mathbf{Q}\right)$ [29]. Let $\sigma_{\max}$ denote the maximum singular value of $R\left(\mathbf{Q}\right)$. Then, the optimal solutions of Problem (53) are given by
$\mathbf{Q}_{\mathrm{v}}$ $=\mathrm{unvec}_{N_{\mathrm{v}}\times J_{\mathrm{v}}}\left(\sqrt{\sigma_{\max}}\mathbf{u}_{\max}\right),$ (55)
$\mathbf{Q}_{\mathrm{h}}$ $=\mathrm{unvec}_{N_{\mathrm{h}}\times J_{\mathrm{h}}}\left(\sqrt{\sigma_{\max}}\mathbf{v}_{\max}^{*}\right),$ (56)
where ${\mathbf{u}}_{\max}$ and ${\mathbf{v}}_{\max}$ represent the left and right singular vectors corresponding to $\sigma_{\max}$, respectively. Thus, (22) can be rewritten as follows
$\quad\mathbf{a}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}},\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)$
$=\left(\mathbf{Q}_{v}^{\mathrm{H}}\otimes\mathbf{Q}_{h}^{\mathrm{H}}\right)\left(\mathbf{a}_{r}^{\mathrm{v}}\left(\varPhi_{k}^{\mathrm{r}}\right)\otimes\mathbf{a}_{r}^{\mathrm{h}}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}}\right)\right)\mathbf{a}_{t}^{\mathrm{H}}\left(\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)\mathbf{F1}$
$=\left(\mathbf{Q}_{v}^{\mathrm{H}}\mathbf{a}_{r}^{\mathrm{v}}\left(\varPhi_{k}^{\mathrm{r}}\right)\right)\otimes\left(\mathbf{Q}_{h}^{\mathrm{H}}\mathbf{a}_{r}^{\mathrm{h}}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}}\right)\right)\mathbf{a}_{t}^{\mathrm{H}}\left(\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)\mathbf{F1}$
$\triangleq\mathbf{o}_{k}\otimes\mathbf{w}_{k},$ (57)
where
$\mathbf{o}_{k}=\mathbf{Q}_{v}^{\mathrm{H}}\mathbf{a}_{r}^{\mathrm{v}}\left(\varPhi_{k}^{\mathrm{r}}\right),$ (58)
$\mathbf{w}_{k}=\mathbf{Q}_{h}^{\mathrm{H}}\mathbf{a}_{r}^{\mathrm{h}}\left(\varTheta_{k}^{\mathrm{r}},\varPhi_{k}^{\mathrm{r}}\right)\mathbf{a}_{t}^{\mathrm{H}}\left(\varTheta_{k}^{\mathrm{t}},\varPhi_{k}^{\mathrm{t}}\right)\mathbf{F1}.$ (59)
Moreover, by defining $\mathbf{O}=\left[\mathbf{o}_{1},\dots,\mathbf{o}_{K+1}\right]$ and $\mathbf{W}=\left[\mathbf{w}_{1},\dots,\mathbf{w}_{K+1}\right]$, we can reformulate factor matrix $\mathbf{A}$ as follows
$\mathbf{A}$ $=[\mathbf{o}_{1}\otimes\mathbf{w}_{1},\cdots,\mathbf{o}_{K+1}\otimes\mathbf{w}_{K+1}]$
$=\mathbf{O}\odot\mathbf{W}.$ (60)
To estimate $\mathbf{O}=\left[\mathbf{o}_{1},\dots,\mathbf{o}_{K+1}\right]$ and $\mathbf{W}=\left[\mathbf{w}_{1},\dots,\mathbf{w}_{K+1}\right]$, we formulate the following problem
$\max_{\mathbf{o}_{k},\mathbf{w}_{k}}\left\|\mathrm{unvec}_{J_{\mathrm{v}}\times J_{\mathrm{h}}}(\hat{\mathbf{a}}_{k})-\mathbf{o}_{k}\mathbf{w}_{k}^{\mathrm{T}}\right\|_{\mathrm{F}}^{2}.$ (61)
Problem (61) can be efficiently solved via the rank-1 truncated SVD. Let $\sigma_{k,\max}$, $\mathbf{u}_{k,\max}$, and $\mathbf{v}_{k,\max}$ denote the maximum singular value, and the corresponding left and right singular vectors, respectively, of $\mathrm{unvec}_{J_{\mathrm{v}}\times J_{\mathrm{h}}}(\hat{\mathbf{a}}_{k})$. We can set the solution of Problem (61) as
$\hat{\mathbf{o}}_{k}=\sigma_{k,\max}\mathbf{u}_{k,\max},\ \hat{\mathbf{w}}_{k}=\mathbf{v}^{*}_{k,\max}.$ (62)
Then, Problem (51) can be decoupled as two subproblems:
$\max_{\varPhi_{k}^{\mathrm{r}}}\frac{\left|\hat{\mathbf{o}}_{k}^{\mathrm{H}}\left(\mathbf{Q}_{\mathrm{v}}^{\mathrm{H}}\mathbf{a}_{r}^{\mathrm{v}}\left(\varPhi_{k}^{\mathrm{r}}\right)\right)\right|^{2}}{\left\|\hat{\mathbf{o}}_{k}\right\|_{2}^{2}\left\|\left(\mathbf{Q}_{\mathrm{v}}^{\mathrm{H}}\mathbf{a}_{r}^{\mathrm{v}}\left(\varPhi_{k}^{\mathrm{r}}\right)\right)\right\|_{2}^{2}},$ (63)
$\max_{\varTheta_{k}^{\mathrm{r}}}\frac{\left|\hat{\mathbf{w}}_{k}^{\mathrm{H}}\left(\mathbf{Q}_{\mathrm{h}}^{\mathrm{H}}\mathbf{a}_{r}^{\mathrm{h}}\left(\varTheta_{k}^{\mathrm{r}},\hat{\varPhi}_{k}^{\mathrm{r}}\right)\right)\right|^{2}}{\left\|\hat{\mathbf{w}}_{k}\right\|_{2}^{2}\left\|\left(\mathbf{Q}_{\mathrm{h}}^{\mathrm{H}}\mathbf{a}_{r}^{\mathrm{h}}\left(\varTheta_{k}^{\mathrm{r}},\hat{\varPhi}_{k}^{\mathrm{r}}\right)\right)\right\|_{2}^{2}}.$ (64)
These subproblems can be solved sequentially through one-dimensional search method. From (7), given the optimal solutions $\hat{\varTheta}_{k}^{\mathrm{r}}$ and $\hat{\varTheta}_{k}^{\mathrm{r}}$ of subproblems (63) and (64), the azimuth and elevation angles of AoA can be calculated as
$\hat{\theta}_{k}^{\mathrm{r}}=\mathrm{arc}\cos\mathrm{(}\hat{\varPhi}_{k}^{\mathrm{r}}),\ \hat{\phi}_{k}^{\mathrm{r}}=\mathrm{arc}\cos\left(\frac{\hat{\varTheta}_{k}^{\mathrm{r}}}{\sin\mathrm{(}\hat{\theta}_{k}^{\mathrm{r}})}\right).$ (65)
In Algorithm 1, we summarize the procedures of the proposed SVD-based algorithm for solving Problem (51) and estimating the AoAs $\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\}_{k=1}^{K+1}$.
Algorithm 1 SVD-based Decoupling AoA Estimation Algorithm
1:Combining matrix $\mathbf{Q}$ and recovered factor matrix $\hat{\mathbf{A}}$.
2:The estimation of azimuth/elevation AoAs $\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\}_{k=1}^{K+1}$.
3:Initialization: Set $J_{\mathrm{v}}$ and $J_{\mathrm{h}}$ subject to $J_{\mathrm{v}}J_{\mathrm{h}}=R$;
4:Calculate $R\left(\mathbf{Q}\right)$ via (54);
5:Obtain $\mathbf{Q}_{\mathrm{v}}$ and $\mathbf{Q}_{\mathrm{h}}$ via (55) and (56), respectively;
6:for $k=1:K+1$ do
7:     Obtain $\hat{\mathbf{o}}_{k}$ and $\hat{\mathbf{w}}_{k}$ via (62);
8:     Obtain $\hat{\varPhi}_{k}^{\mathrm{r}}$ and $\hat{\varTheta}_{k}^{\mathrm{r}}$ by solving Problem (63) and Problem (64) via the one-dimensional search, respectively;
9:     Calculate $\hat{\theta}_{k}^{\mathrm{r}}$ and $\hat{\phi}_{k}^{\mathrm{r}}$ via (65);
10:end for
III-D Uniqueness Analysis
This subsection establishes uniqueness guarantees for the tensor decomposition (25) to ensure reliable recovery of factor matrices and subsequent parameter estimation. By extending the classical Kruskal’s uniqueness theorem [30] and incorporating structural constraints inherent to the factor matrices, we have the following relaxed uniqueness criterion [21, 31]:
Lemma 1
Consider a third-order tensor $\mathcal{X}\in\mathbb{C}^{I_{1}\times I_{2}\times I_{3}}$ with rank $K$ decomposed into factor matrices $\mathbf{A}^{(1)}\in\mathbb{C}^{I_{1}\times K}$, $\mathbf{A}^{(2)}\in\mathbb{C}^{I_{2}\times K}$ and $\mathbf{A}^{(3)}\in\mathbb{C}^{I_{3}\times K}$, where $\mathbf{A}^{(3)}$ is a Vandermonde matrix with generators $\{z_{k}\}_{k=1}^{K}$. Let $k(\mathbf{A})$ denote the Kruskal-rank of matrix $\mathbf{A}$. The decomposition achieves uniqueness if the following joint rank conditions hold:
$k\left(\mathbf{A}^{(L_{1},3)}\odot\mathbf{A}^{(2)}\right)=K,$ (66)
$k\left(\mathbf{A}^{(L_{2},3)}\odot\mathbf{A}^{(1)}\right)=K,$ (67)
where $\mathbf{A}^{(L_{1},3)}$ represents the first $L_{1}$ rows of $\mathbf{A}^{(3)}$, and $\mathbf{A}^{(L_{2},3)}$ denotes the first $L_{2}$ rows of $\mathbf{A}^{(2)}$. The uniqueness means that any possible combination of factor matrices $(\hat{\mathbf{A}}^{(1)},\mathrm{}\hat{\mathbf{A}}^{(2)},\hat{\mathbf{A}}^{(3)})$ for $\mathcal{X}$ satisfies
$\hat{\mathbf{A}}^{(1)}=\mathbf{A}^{(1)}\mathbf{\Pi\Delta}_{1},\,\mathrm{}\hat{\mathbf{A}}^{(2)}=\mathbf{A}^{(2)}\mathbf{\Pi\Delta}_{2},\,\hat{\mathbf{A}}^{(3)}=\mathbf{A}^{(3)}\mathbf{\Pi\Delta}_{3},$ (68)
where $\mathbf{\Pi}$ is a $K\times K$ permutation matrix, and $\mathbf{\Delta}_{1}$, $\mathbf{\Delta}_{2}$, and $\mathbf{\Delta}_{3}$ with $\mathbf{\Delta}_{1}\mathbf{\Delta}_{2}\mathbf{\Delta}_{3}=\mathbf{I}_{K}$ are diagonal matrices. For generic scenarios, these conditions reduce to the simplified dimensionality constraint:
$\min\left(\left(L_{1}-1\right)I_{2},\,L_{2}I_{1}\right)\geqslant K.$ (69)
[…]
In the considered system model, the tensor dimensions $I_{1}$, $I_{2}$ and $I_{3}$ correspond respectively to the number of RF chains $R$, OFDM symbols $N$ and subcarriers $M$. The rank of tensor $\bm{\mathcal{Y}}$ is $K+1$. Therefore, sufficient conditions for the decomposition uniqueness can be formulated as follows
$L_{1}+L_{2}$ $=M+1,$ (70)
$(L_{1}-1)N$ $\geqslant K+1,$ (71)
$L_{2}R$ $\geqslant K+1.$ (72)
These constraints can be satisfied by appropriately selecting parameters, ensuring automatic association of the extracted parameters with respective propagation paths. Note that telecommunication systems in 5G and beyond typically employ a large number of subcarriers $M$. This enables flexible selection of integers $L_{1}$ and $L_{2}$.
III-E Algorithm Development
Algorithm 2 Tensor-Based Sensing Parameter Estimation Algorithm
1:Received signal tensor $\bm{\mathcal{Y}}$.
2:Estimations of bistatic ranges, bistatic Doppler velocities, and azimuth/elevation angles of arrival $\left\{d_{k},v_{k},\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\right\}_{k=1,k\neq k_{\mathrm{LoS}}}^{K+1}$.
3:Initialization: Mode-1 unfold tensor $\bm{\mathcal{Y}}$ as $\mathbf{Y}_{(1)}$ in (27); Set smoothing parameters $L_{1}$ and $L_{2}$ subject to $L_{1}+L_{2}=M+1$;
4:Calculate the smoothed measurement matrix $\mathbf{Y}_{\mathrm{S}}$ via (III-B);
5:Obtain $\mathbf{U}$, $\mathbf{\Sigma}$, and $\mathbf{V}$ by performing SVD on $\mathbf{Y}_{\mathrm{S}}$ in (30);
6:Obtain $\hat{\mathbf{Z}}$ and $\hat{\mathbf{M}}$ by performing EVD on $\mathbf{U}_{1}^{\dagger}\mathbf{U}_{2}$ in (36);
7:Construct columns of $\hat{\mathbf{C}}$ via (37);
8:Construct columns of $\hat{\mathbf{B}}$ via (39);
9:Construct columns of $\hat{\mathbf{A}}$ via (40);
10:Obtain the index of LoS path $k_{\mathrm{LoS}}$ via (43);
11:for $k\in\{1,\cdots,K+1\}\setminus k_{\mathrm{LoS}}$ do
12:     Calculate $\hat{d}_{k}$ via (46);
13:     Calculate $\hat{v}_{k}$ via (50);
14:     Obtain $\hat{\theta}_{k}$ and $\hat{\phi}_{k}$ following Step 7-9 of Algorithm 1;
15:end for
Based on the discussions in Sections III-B and III-C, we summarize the proposed tensor decomposition and spatial smoothing-based algorithm for sensing parameter estimation in Algorithm 2.
Complexity Analysis: The main computational complexity is analyzed as follows:
  • Step 2: The computational complexity of constructing $\mathbf{Y}_{\mathrm{S}}$ is initially $\mathcal{O}(L_{1}L_{2}NR)$. However, by leveraging the sparse cyclic selection matrices $\mathbf{J}_{l}$ to extract and assemble specific rows from $\mathbf{Y}_{(1)}$ into the extended matrix formulation, this complexity can be reduced to practically negligible levels.
  • Step 3: The complexity of performing truncated SVD on $\mathbf{Y}_{S}$ is $\mathcal{O}(L_{1}L_{2}NRK)$.
  • Step 4: Calculating the pseudo-inverse of $\mathbf{U}_{1}$ has a complexity of $\mathcal{O}(L_{1}NK)$. Besides, performing EVD on $\mathbf{U}_{1}^{\dagger}\mathbf{U}_{2}$ has a complexity of $\mathcal{O}(K^{3})$.
  • Step 5-7: The construction of matrices $\mathbf{C}$, $\mathbf{B}$, and $\mathbf{A}$ primarily requires the complexity of $\mathcal{O}(KM)$, $\mathcal{O}(L_{1}N^{2}K)$, and $\mathcal{O}(L_{2}R^{2}K)$, respectively.
  • Step 9-13: Denote by $G$ the number of iterations in the one-dimensional search method. In the parameter extraction stage, the complexity is dominated by that of estimating $\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\}_{k=1,k\neq k_{\mathrm{LoS}}}^{K+1}$, which is given by $\mathcal{O}(RKG(J_{\mathrm{v}}N_{\mathrm{v}}+J_{\mathrm{h}}N_{\mathrm{h}}))$.
In general, the total complexity of Algorithm 2 can be summarized as $\mathcal{O}(L_{1}L_{2}NRK+L_{1}N^{2}K+L_{2}R^{2}K+RKG(J_{\mathrm{v}}N_{\mathrm{v}}+J_{\mathrm{h}}N_{\mathrm{h}}))$.
Remark 2
In the previous analysis, $k\in\left\{1,\dots,K+1\right\}$ has been used to index the propagation paths between tBS $n_{\mathrm{t}}$ and rBS $n_{\mathrm{r}}$. Since the index $k_{\mathrm{LoS}}$ can be obtained at step 10 of Algorithm 2, we reallocate the index domain $k\in\left\{1,\dots,K\right\}$ to identify the $K$ targets in the following formulations. In addition, we will preserve the subscripts $n_{\mathrm{t}}$ and $n_{\mathrm{r}}$ in all sensing parameters to highlight the association with their respective tBS-rBS pairs, departing from the simplified notations in the previous discussions.
IV Position and Velocity Estimation
Based on the bistatic ranges $\{\hat{d}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}\}$ and AoA parameters $\{\hat{\theta}_{n_{\mathrm{t}},k,n_{\mathrm{r}}},\hat{\phi}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}\}$ estimated from a distributed tBS-rBS pair, the system can theoretically estimate the 3D positions of all targets. However, since only bistatic Doppler velocities $\{\hat{v}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}\}$ are measurable at each rBS, resolving the true 3D velocities necessitates fusing at least three independent radial velocity estimates. To tackle this problem and improve the accuracy of position and velocity estimation, this section proposes a MST-based cooperative data fusion algorithm. This scheme jointly resolves the data association ambiguity among estimated parameters from different tBS-rBS pairs while jointly optimizing the 3D positions and velocities of all targets.
IV-A Basic Target Localization
Let $d_{n_{\mathrm{t}},k}^{\mathrm{t}}=\left\|\mathbf{p}_{k}^{\mathrm{u}}-\mathbf{p}_{n_{\mathrm{t}}}^{\mathrm{t}}\right\|_{2}$ denote the distance between tBS $n_{\mathrm{t}}$ and target $k$, where $\mathbf{p}_{k}^{\mathrm{u}}=\left[x_{k}^{\mathrm{u}},y_{k}^{\mathrm{u}},z_{k}^{\mathrm{u}}\right]^{\mathrm{T}}$ and $\mathbf{p}_{n_{\mathrm{t}}}^{\mathrm{t}}=\left[x_{n_{\mathrm{t}}}^{\mathrm{t}},y_{n_{\mathrm{t}}}^{\mathrm{t}},z_{n_{\mathrm{t}}}^{\mathrm{t}}\right]^{\mathrm{T}}$ are the 3D position vectors of tBS $n_{\mathrm{t}}$ and target $k$, respectively. Similarly, the distance between target $k$ and rBS $n_{\mathrm{r}}$ is defined as $d_{k,n_{\mathrm{r}}}^{\mathrm{r}}=\left\|\mathbf{p}_{k}^{\mathrm{u}}-\mathbf{p}_{n_{\mathrm{r}}}^{\mathrm{r}}\right\|_{2}$, where $\mathbf{p}_{n_{\mathrm{r}}}^{\mathrm{r}}=\left[x_{n_{\mathrm{r}}}^{\mathrm{r}},y_{n_{\mathrm{r}}}^{\mathrm{r}},z_{n_{\mathrm{r}}}^{\mathrm{r}}\right]^{\mathrm{T}}$ represents the position of rBS $n_{\mathrm{r}}$. Given $\hat{d}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$, $\hat{\theta}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$, and $\hat{\phi}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$, the 3D position $\mathbf{p}_{k}^{\mathrm{u}}$ of target $k$ can be determined by solving the following system of equations:
$\begin{cases}d_{n_{\mathrm{t}},k}^{\mathrm{t}}+d_{k,n_{\mathrm{r}}}^{\mathrm{r}}=\hat{d}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}\\ \mathbf{p}_{k}^{\mathrm{u}}=\mathbf{p}_{n_{\mathrm{r}}}^{\mathrm{t}}+d_{k,n_{\mathrm{r}}}^{\mathrm{r}}\hat{\mathbf{r}}^{\mathrm{r}}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}\\ \end{cases},$ (73)
where $\hat{\mathbf{r}}^{\mathrm{r}}_{n_{\mathrm{r}},k,n_{\mathrm{r}}}$ represents the unit direction vector from target $k$ to rBS $n_{\mathrm{r}}$ derived from AoA estimations, expressed in the global coordinate system as:
$\hat{\mathbf{r}}^{\mathrm{r}}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}=\mathbf{T}\left(\chi_{n_{\mathrm{r}}}^{\mathrm{r}}\right)\left[\begin{array}[]{c}\sin\hat{\theta}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{r}}\cos\hat{\phi}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{r}}\\ \sin\hat{\theta}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{r}}\sin\hat{\phi}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{r}}\\ \cos\hat{\phi}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{r}}\\ \end{array}\right],$ (77)
$\mathbf{T}\left(\chi_{n_{\mathrm{r}}}^{\mathrm{r}}\right)=\left[\begin{matrix}\cos\chi_{n_{\mathrm{r}}}^{\mathrm{r}}&\sin\chi_{n_{\mathrm{r}}}^{\mathrm{r}}&0\\ -\sin\chi_{n_{\mathrm{r}}}^{\mathrm{r}}&\cos\chi_{n_{\mathrm{r}}}^{\mathrm{r}}&0\\ 0&0&1\\ \end{matrix}\right],$ (78)
where $\chi_{n_{\mathrm{r}}}^{\mathrm{r}}$ is the horizontal orientation of rBS $n_{\mathrm{r}}$’s array, defined as the angle between the array normal and the global $x$-axis. Solving (73) for $\forall k$ yields basic position estimates $\{\hat{\mathbf{p}}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{u}}\}_{k=1}^{K}$ related to tBS $n_{\mathrm{t}}$ and rBS $n_{\mathrm{r}}$. However, independent processing across the tBS-rBS pairs introduces data association ambiguity: due to inconsistent indexing across pairs, the estimated $\hat{\mathbf{p}}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{u}}$ and $\hat{\mathbf{p}}_{n^{\prime}_{\mathrm{t}},k,n^{\prime}_{\mathrm{r}}}^{\mathrm{u}}$ may correspond to different targets when $n_{\mathrm{t}}\neq n^{\prime}_{\mathrm{t}}$ or $n_{\mathrm{r}}\neq n^{\prime}_{\mathrm{r}}$. In addition, estimation errors in $\hat{d}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$, $\hat{\theta}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$, and $\hat{\phi}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$ lead to biases in position estimates. Although greedy algorithms can be applied for the target-to-estimate association [32], the erroneous position estimates cannot be effectively eliminated. This limitation often leads to incorrect data associations or significant global positioning biases, particularly in scenarios with a large number of targets or low signal-to-noise ratios (SNRs). In the following, an effective error elimination and data association algorithm is proposed based on the MST method.
IV-B Error Elimination and Data Association
To establish the MST-based data association framework, we first define an undirected weighted graph $\mathcal{G}=\left(\mathcal{V},\mathcal{E},W\right)$ consisting of the vertex set $\mathcal{V}$, the edge set $\mathcal{E}$, and the weight function $W$ as illustrated in Fig. 3(a). Specifically, $\mathcal{V}$ is represented as a union of $N_{\mathrm{t}}\times N_{\mathrm{r}}$ subsets as follows
$\mathcal{V}=\bigcup_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}{\bigcup_{n_{\mathrm{r}}=1}^{N_{\mathrm{r}}}{\mathcal{V}_{n_{\mathrm{t}},n_{\mathrm{r}}}}},$ (79)
where each subset $\mathcal{V}_{n_{\mathrm{t}},n_{\mathrm{r}}}$ corresponds to the estimates from tBS $n_{\mathrm{t}}$ and rBS $n_{\mathrm{r}}$, indexed as
$\mathcal{V}_{n_{\mathrm{t}},n_{\mathrm{r}}}=\{((n_{\mathrm{r}}-1)N_{\mathrm{t}}+n_{\mathrm{t}}-1)K+1,\dots,((n_{\mathrm{r}}-1)N_{\mathrm{t}}+n_{\mathrm{t}})K\}.$ (80)
The edge $\mathcal{E}$ connects vertices across different tBS-rBS pairs, defined as
=$\mathcal{E}=\left\{\left(e_{1},e_{2}\right)|e_{1}\in V_{n_{\mathrm{t}},n_{\mathrm{r}}},\,e_{2}\in V\backslash V_{n_{\mathrm{t}},n_{\mathrm{r}}}\right\}.$ (81)
The edge weight between any two vertices $e_{1}$ and $e_{2}$ is defined as the Euclidean distance between their corresponding position estimates as follows
$W\left(e_{1},e_{2}\right)=\left\|\hat{\mathbf{p}}_{\xi^{\mathrm{t}}\left(e_{1}\right),\xi^{\mathrm{u}}\left(e_{1}\right),\xi^{\mathrm{r}}\left(e_{1}\right)}-\hat{\mathbf{p}}_{\xi^{\mathrm{t}}\left(e_{2}\right),\xi^{\mathrm{u}}\left(e_{2}\right),\xi^{\mathrm{r}}\left(e_{2}\right)}\right\|_{2},$
$\forall\left(e_{1},e_{2}\right)\in\mathcal{E},$ (82)
[…]
It can be noted that valid position estimates of the same target should cluster spatially, while erroneous estimates deviate significantly. That is, a vertex is likely corresponding to erroneous estimate if all its adjacent edges have large weights. To eliminate such outliers, we introduce a threshold $\varpi$ and define the following pruned edge set
=$\mathcal{E}^{\prime}=\left\{\left(e_{1},e_{2}\right)\in\mathcal{E}|W_{\min}\left(e_{1}\right)>\varpi\right\},$ (83)
where $W_{\min}\left(e_{1}\right)$ denotes the minimum weight among all edges adjacent to $e_{1}$. By removing these edges, the graph $\mathcal{G}$ is updated as
$\tilde{\mathcal{G}}=\left(\mathcal{V},\mathcal{E}\backslash\mathcal{E}^{\prime},W\right),$ (84)
which isolates vertices corresponding to erroneous estimates as illustrated in Fig. 3b.
Finally, the MST of the pruned graph $\tilde{\mathcal{G}}$ can be computed using the well-established algorithms such as Prim’s [33] or Kruskal’s [34]. Since the valid position estimates are clustered, the edges connecting estimates of the same target usually exhibit smaller weights and are preferentially retained in the MST. Therefore, the final data association is achieved by removing the $K-1$ longest edges from the MST. This operation partitions the MST into $K$ sub-graphs as shown in Fig. 3(c), each corresponding to a unique target.
Figure 3: An illustration of the MST-based error elimination and data association process, where the numbers of targets and tBS-rBS pairs are $K=3$ and $N_{\mathrm{t}}N_{\mathrm{r}}=3$, receptively.
IV-C Data Fusion for Target Localization
Based on the MST-based association results in the previous subsection, we develop a data fusion scheme for cooperative target localization. While the subsequent derivations retain summation over all sensing parameters for notational continuity, we emphasize that only the validated parameters (after error elimination) contribute to the final estimation.
The fused position estimates $\left\{\accentset{*}{\mathbf{p}}_{k}^{\mathrm{u}}\right\}_{k=1}^{K}$ can be obtained through either hard or soft fusion [16]. In the hard fusion approach, basic estimates $\{\hat{\mathbf{p}}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{u}}\}$ are directly combined, for instance via arithmetic averaging:
$\accentset{*}{\mathbf{p}}_{k}^{\mathrm{u},\mathrm{ave}}=\frac{1}{N_{\mathrm{t}}N_{\mathrm{r}}}\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}{\sum_{n_{\mathrm{r}}=1}^{N_{\mathrm{r}}}{\hat{\mathbf{p}}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{u}}}}.$ (85)
To achieve higher localization accuracy, we adopt a soft fusion paradigm by formulating a maximum-likelihood estimator that jointly processes all original sensing parameters $\{\hat{d}_{n_{\mathrm{t}},k,n_{\mathrm{r}}},\hat{\theta}_{n_{\mathrm{t}},k,n_{\mathrm{r}}},\hat{\phi}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}\}$. Specifically, we define the following loss function
$f\left(\mathbf{p}_{k}^{\mathrm{u}}\right)=$ $\frac{\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}{\sum_{n_{\mathrm{r}}=1}^{N_{\mathrm{r}}}{\alpha_{n_{\mathrm{t}},k,n_{\mathrm{r}}}\left|\hat{d}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}-d_{n_{\mathrm{t}},k}^{\mathrm{t}}-d_{k,n_{\mathrm{r}}}^{\mathrm{r}}\right|}}}{\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}{\sum_{n_{\mathrm{r}}=1}^{N_{\mathrm{r}}}{\alpha_{n_{\mathrm{t}},k,n_{\mathrm{r}}}}}}$
$+\frac{\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}{\sum_{n_{\mathrm{r}}=1}^{N_{\mathrm{r}}}{\beta_{n_{\mathrm{t}},k,n_{\mathrm{r}}}\left\|\hat{\mathbf{r}}^{\mathrm{r}}_{n_{\mathrm{r}},k,n_{\mathrm{r}}}-\mathbf{r}^{\mathrm{r}}_{k,n_{\mathrm{r}}}\right\|}}}{\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}{\sum_{n_{\mathrm{r}}=1}^{N_{\mathrm{r}}}{\beta_{n_{\mathrm{t}},k,n_{\mathrm{r}}}}}},$ (86)
where $\mathbf{r}_{k,n_{\mathtt{r}}}^{\mathtt{r}}=\left(\mathbf{p}_{n_{\mathtt{r}}}^{\mathtt{r}}-\mathbf{p}_{k}^{\mathtt{u}}\right)/\left\|\mathbf{p}_{n_{\mathtt{r}}}^{\mathtt{r}}-\mathbf{p}_{k}^{\mathtt{u}}\right\|_{2}$ denotes the unit direction vector from target $k$ to rBS $n_{\mathrm{r}}$ determined by the target position $\mathbf{p}_{k}^{\mathrm{u}}$. The weights $\alpha_{n_{\mathrm{t}},k,n_{\mathrm{r}}}=(d_{n_{\mathrm{t}},k}^{\mathrm{t}}d_{k,n_{\mathrm{r}}}^{\mathrm{r}})^{-2}$ and $\beta_{n_{\mathrm{t}},k,n_{\mathrm{r}}}=\alpha_{n_{\mathrm{t}},k,n_{\mathrm{r}}}d_{k,n_{\mathrm{t}}}^{\mathrm{r}}$ are designed based on the fact that the signal path loss is inversely proportional to $\left(d_{n_{\mathrm{t}}}^{\mathrm{t}}\left(\mathbf{p}_{k}^{\mathrm{u}}\right)\cdot d_{n_{\mathrm{r}}}^{\mathrm{r}}\left(\mathbf{p}_{k}^{\mathrm{u}}\right)\right)^{2}$, thereby assigning higher confidence to sensing parameters derived under stronger SNRs [16]. Then, the optimal position estimate $\accentset{*}{\mathbf{p}}_{k}^{\mathrm{u}}$ is obtained by solving the following problem
$\min_{\mathbf{p}_{k}^{\mathrm{u}}}f\left(\mathbf{p}_{k}^{\mathrm{u}}\right).$ (87)
Problem (87) is an unconstrained optimization problem, which can be efficiently solved by the quasi-Newton algorithm [35]. In this work, we initialize the quasi-Newton solver with the averaging-based estimate $\accentset{*}{\mathbf{p}}_{k}^{\mathrm{u},\mathrm{ave}}$, leveraging its proximity to the true position to mitigate the sensitivity of non-convex optimization to initial conditions.
IV-D Velocity Estimation
Leveraging the cooperative position estimates obtained in the previous subsection, we estimate the true 3D velocity $\mathbf{v}_{k}=\left[v_{x},v_{y},v_{z}\right]^{\mathrm{T}}$ of target $k$ using its bistatic Doppler velocity estimates $\{\hat{v}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}|\forall n_{\mathrm{t}},n_{\mathrm{r}}\}$. Here, the sensing parameters $d_{n_{\mathrm{t}},k}^{\mathrm{t}}$, $d_{k,n_{\mathrm{r}}}^{\mathrm{r}}$, $\mathbf{r}_{k,n_{\mathrm{r}}}^{\mathrm{r}}$, and $\alpha_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$ are recalculated using $\accentset{*}{\mathbf{p}}_{k}^{\mathrm{u}}$ instead of the unknown true position $\mathbf{p}_{k}^{\mathrm{u}}$ , with $\mathbf{r}_{n_{\mathrm{t}},k}^{\mathrm{t}}$ denoting the unit direction vector from target $k$ to tBS $n_{\mathrm{t}}$.
The velocity estimation can be formulated as the following weighted least squares problem
$\min_{\mathbf{v}_{k}}\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}{\sum_{n_{\mathrm{r}}=1}^{N_{\mathrm{r}}}{w_{n_{\mathrm{t}},k,n_{\mathrm{r}}}\left(\hat{v}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}-\left(\mathbf{r}_{n_{\mathrm{t}},k}^{\mathrm{t}}+\mathbf{r}_{k,n_{\mathrm{r}}}^{\mathrm{r}}\right)^{\mathrm{T}}\mathbf{v}_{k}\right)^{2},}}$ (88)
where the weights $w_{n_{\mathrm{t}},k,n_{\mathrm{r}}}$ inherit the path-loss-aware scaling from the position estimation:
$w_{n_{\mathrm{t}},k,n_{\mathrm{r}}}=\frac{\alpha_{n_{\mathrm{t}},k,n_{\mathrm{r}}}}{\sum_{n_{\mathrm{t}}=1}^{N_{\mathrm{t}}}{\sum_{n_{\mathrm{r}}=1}^{N_{\mathrm{r}}}{\alpha_{n_{\mathrm{t}},k,n_{\mathrm{r}}}}}}.$ (89)
To solve this problem efficiently, we vectorize it as
$\min_{\mathbf{v}_{k}}\left(\hat{\mathbf{v}}_{k}-\hat{\mathbf{R}}_{k}\mathbf{v}_{k}\right)^{\mathrm{T}}\mathbf{W}_{k}\left(\hat{\mathbf{v}}_{k}-\hat{\mathbf{R}}_{k}\mathbf{v}_{k}\right),$ (90)
where $\hat{\mathbf{v}}_{k}\in\mathbb{R}^{N_{\mathrm{t}}N_{\mathrm{r}}\times 1}$, $\hat{\mathbf{R}}_{k}\in\mathbb{R}^{N_{\mathrm{t}}N_{\mathrm{r}}\times 3}$, and $\mathbf{W}_{k}\in\mathbb{R}^{N_{\mathrm{t}}N_{\mathrm{r}}\times N_{\mathrm{t}}N_{\mathrm{r}}}$ are constructed as
$\hat{\mathbf{v}}_{k}=\left[\hat{v}_{1,k,1},\dots,\hat{v}_{N_{\mathrm{t}}k,N_{\mathrm{r}}}\right]^{\mathrm{T}},$ (91)
$\hat{\mathbf{R}}_{k}=\left[\left(\mathbf{r}_{1,k}^{\mathrm{t}}+\mathbf{r}_{k,1}^{\mathrm{r}}\right),\dots,\left(\mathbf{r}_{N_{\mathrm{t}},k}^{\mathrm{t}}+\mathbf{r}_{k,N_{\mathrm{r}}}^{\mathrm{r}}\right)\right]^{\mathrm{T}},$ (92)
$\mathbf{W}_{k}=\mathrm{diag}\left(\left[w_{1,k,1},\dots,w_{N_{\mathrm{t}},k,N_{\mathrm{r}}}\right]\right).$ (93)
Recall that $N_{\mathrm{t}}N_{\mathrm{r}}\geqslant 3$ independent radial velocity estimates are required for 3D velocity estimation, which means $\mathrm{rank}(\hat{\mathbf{R}}_{k})=3$. Hence, the closed-form optimal solution of Problem (90) is given by
$\accentset{*}{\mathbf{v}}_{k}=\left(\hat{\mathbf{R}}_{k}^{\mathrm{T}}\mathbf{W}_{k}\hat{\mathbf{R}}_{k}\right)^{-1}\hat{\mathbf{R}}_{k}^{\mathrm{T}}\mathbf{W}_{k}\hat{\mathbf{v}}_{k}.$ (94)
IV-E Algorithm Development
Algorithm 3 Cooperative Position and Velocity Estimation Algorithm
1:Sensing parameters $\{\hat{d}_{n_{\mathrm{t}},k,n_{\mathrm{r}}},\hat{\theta}_{n_{\mathrm{t}},k,n_{\mathrm{r}}},\hat{\phi}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}\hat{v}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}\}$.
2:3D position and velocity estimates of targets $\{\accentset{*}{\mathbf{p}}_{k}^{\mathrm{u}},\accentset{*}{\mathbf{v}}_{k}\}$.
3:for $\forall n_{\mathrm{t}},k,n_{\mathrm{r}}$ do $\triangleright$ Basic position estimation
4:     Obtain $\hat{\mathbf{p}}_{n_{\mathrm{t}},k,n_{\mathrm{r}}}^{\mathrm{u}}$ by solving system of equations (73);
5:end for
6:Construct undirected weighted graph $\mathcal{G}=\left(\mathcal{V},\mathcal{E},W\right)$ via (79)-(IV-B); $\triangleright$ Data association
7:Obtain graph $\tilde{\mathcal{G}}$ via (84);
8:Generate the MST of $\tilde{\mathcal{G}}$;
9:Obtain sub-graphs by removing the $K-1$ longest edges from the MST;
10:Obtain data association results from vertices in each subgraph;
11:for $\forall k$ do $\triangleright$ Cooperative position estimation
12:     Obtain $\accentset{*}{\mathbf{p}}_{k}^{\mathrm{u}}$ by solving Problem (87);
13:end for
14:for $\forall k$ do $\triangleright$ Velocity estimation
15:     Calculate $\accentset{*}{\mathbf{v}}_{k}$ via (94);
16:end for
Based on the above discussions, we summarize the proposed cooperative position and velocity estimation algorithm in Algorithm 3. The algorithm comprises four cascaded phases: 1) basic position estimation, 2) MST-based data association, 3) cooperative position optimization, and 4) velocity estimation.
Complexity Analysis: The computational complexity of Algorithm 3 is dominated by the MST generation in Step 8 and solving Problem (87) in Step 12. The MST generation via Kruskal’s algorithm incurs a complexity of $𝑂\superscript\superscriptO\left(\left|\mathcal{E}\backslash\mathcal{E}^{\prime}\right|\log\left|\mathcal{E}\backslash\mathcal{E}^{\prime}\right|\right)$, where |\$\left|\mathcal{E}\backslash\mathcal{E}^{\prime}\right|\leqslant\left|\mathcal{E}\right|<\left(KN_{\mathrm{t}}N_{\mathrm{r}}\right)^{2}$. Given typical ISAC network scales, this simplifies to $O(\left(KN_{\mathrm{t}}N_{\mathrm{r}}\right)^{2}\log\left(KN_{\mathrm{t}}N_{\mathrm{r}}\right))$. To solve Problem (87), the computational complexity of each quasi-Newton iteration is of $O\left(N_{\mathrm{t}}N_{\mathrm{r}}\right)$. With the averaging-based initialization $\accentset{*}{\mathbf{p}}_{k}^{\mathrm{u}}$ in (85), the required iteration number $G^{\prime}$ remains small (typically $G^{\prime}\leqslant 5$). Therefore, the total complexity of generating the fused position estimates $\left\{\accentset{*}{\mathbf{p}}_{k}^{\mathrm{u}}\right\}_{k=1}^{K}$ is of $O\left(KG^{\prime}N_{\mathrm{t}}N_{\mathrm{r}}\right)$. To sum up, the total complexity of Algorithm 3 is given by $O((KN_{\mathrm{t}}N_{\mathrm{r}})^{2}\log\left(KN_{\mathrm{t}}N_{\mathrm{r}}\right)+KG^{\prime}N_{\mathrm{t}}N_{\mathrm{r}})$.
Notably, the computational overhead of data fusion for localization and velocity estimation is negligible compared to the parameter extraction phase, which ensures the scalability of cooperative ISAC systems even with increasing network density.
V Simulation Results
In this section, numerical examples are presented to investigate the performance of the proposed MIMO-OFDM cellular network-based cooperative bistatic sensing framework.
V-A Simulation Setups
The simulation considers a cooperative ISAC system comprising 8 BSs uniformly distributed on a circle with a radius of 500 meters (m), as illustrated in Fig. 4. Each BS is positioned at a height of 30 m and equipped with a uniform planar array (UPA) consisting of 16 horizontal and 24 vertical antenna elements, driven by 64 radio frequency (RF) chains. The horizontal orientation $\chi_{n_{\mathrm{r}}}^{\mathrm{r}}$ of each UPA is configured to point towards the center of the circle. To ensure the uniqueness of the CP decomposition, the elements of the precoding matrix $\mathbf{F}$ and the combining matrix $\mathbf{Q}$ are randomly selected from points on the complex unit circle [19].
Figure 4: Illustration of the system model in simulation.
UAVs are modelled as sensing targets, generated uniformly at random within a circular area of radius 400 m and at heights ranging between 50 m and 300 m. The UAVs exhibit a maximum flight speed of 60 km/h and a minimum speed of 5 km/h, with a safety separation distance of 10 m enforced to prevent collisions. The radar cross-section (RCS) of each UAV is set to $\sigma_{\mathrm{RCS}}=0.1$ m2.
Sensing signal transmission utilizes a typical sub-6 GHz configuration with a carrier frequency of $f_{\mathrm{c}}=4.9$ GHz and a subcarrier spacing of $\Delta f=30$ kHz. Unless otherwise stated, each tBS occupies a BWP of 20 MHz (including guard intervals), transmitting sensing signals over 51 resource blocks (RBs), where each RB comprises 12 subcarriers [36]. Each sensing task employs $N=7$ consecutive OFDM symbols. The path losses for the target-scattered paths and the LoS paths are respectively modelled as follows [37]
$\mathrm{PL}_{\mathrm{Ts}}=103.4+20\lg f_{\mathrm{c}}+20\lg d^{\mathrm{t}}+20\lg d^{\mathrm{r}}-10\lg\sigma_{\mathrm{RCS}},$ (95)
$\mathrm{PL}_{\mathrm{LoS}}=32.4+20\lg f_{\mathrm{c}}+20\lg d^{\mathrm{LoS}},$ (96)
where $d^{\mathrm{t}}$ and $d^{\mathrm{r}}$ are the distances in km from the target to the tBS and rBS, respectively, and $d^{\mathrm{LoS}}$ denotes the distance in km between the tBS and rBS.
The root mean square error (RMSE) is employed to quantify the sensing accuracy in the simulations. For any sensing quantity $\mathbf{x}$, which may represent bistatic ranges, Doppler velocities, AoA estimates, or 3D position/velocity estimates, the RMSE is defined as:
$\mathrm{RMSE}\left(\mathbf{x}\right)=\sqrt{\frac{1}{K}\sum_{k=1}^{K}{\left\|\hat{\mathbf{x}}_{k}-\mathbf{x}_{k}\right\|_{2}^{2}}},$ (97)
where $\hat{\mathbf{x}}_{k}$ and $\mathbf{x}_{k}$ denote the estimated and true value for target $k$, respectively. All simulation results are derived from 500 independent Monte Carlo trials. To mitigate the influence of outliers, the reported results represent the average over the top 95% of samples with the highest estimation accuracy, excluding the bottom 5%. Note that all simulation results are presented on a logarithmic scale for the y-axis.
V-B Sensing Parameter Extraction Performance
Recall that all rBSs in the proposed sensing framework execute identical parameter extraction algorithms. To evaluate the parameter extraction performance of the proposed scheme, we first consider a bistatic sensing scenario with a single tBS-rBS pair deployed at (-500 m, 0) and (500 m, 0). The core challenge in estimating bistatic ranges $\left\{d_{k}\right\}$, azimuth/elevation angles $\left\{\theta_{k}^{\mathrm{r}},\phi_{k}^{\mathrm{r}}\right\}$, and bistatic Doppler velocities $\left\{v_{k}\right\}$ is to recover the factor matrices $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{C}$ by solving Problem (26). For performance benchmarking, we introduce a baseline scheme termed ALS. This approach solves Problem (26) using the ALS algorithm [19, 24], serving as a comparative reference to the proposed Algorithm 2.
Figure 5: CPU time versus the number of targets $K$.
Fig. 5 compares the computational complexity of the proposed scheme and the ALS baseline, measured by CPU runtime versus the number of sensing targets $K$. The proposed algorithm demonstrates excellent computational efficiency, with its runtime showing almost negligible growth when the number of targets exceeds $K=3$. On the contrary, the ALS scheme requires higher computational resources in all scenarios, with its runtime increasing approximately exponentially as $K$ grows. Moreover, the proposed algorithm’s runtime at $K=6$ targets remains lower than ALS’s at $K=2$, highlighting its scalability advantage. The computational efficiency of this algorithm, which is unaffected by the growth in the number of targets, enables its practical deployment in real-time multi-target sensing applications, particularly in dense environments where system implementation is constrained by computational resources.
(a) Bistatic range
(b) Azimuth angle
(c) Elevation angle
(d) Bistatic Doppler velocity
Figure 6: Parameter estimation accuracy versus transmit power.
Fig. 6 presents the parameter estimation RMSE performance for $K=\left\{2,3\right\}$ targets under varying transmit power at the tBS. It can be observed that both schemes show monotonic estimation accuracy improvement with increased tBS transmit power, resulting from the increment of SNR. For all sensing parameters, the proposed algorithm demonstrates significant advantages in accuracy, particularly in bistatic range extraction, where the advantage reaches one to two orders of magnitude, achieving millimetre-level precision under sufficient transmit power. Conversely, the ALS scheme suffers from slow and unstable convergence, with limited estimation accuracy. Based on these parameter estimates, Fig. 7 evaluates the RMSE of the basic target position estimation derived from (73). Thanks to the high-precision parameter extraction ability of Algorithm 2, the proposed scheme achieves a localization accuracy approximately one order of magnitude higher than the ALS scheme, with a decimetre-level positioning accuracy when the transmit power is higher than 55 dBm.
Figure 7: Basic localization accuracy versus transmit power.
V-C Performance for 3D Position and Velocity Estimation
In the following, we evaluate the 3D position and velocity estimation performance of the proposed cooperative bistatic ISAC framework with Algorithm 2 and 3. In each generation of the simulation, $N_{t}=2$ tBSs are randomly selected from the 8 available BSs. Then, $N_{\mathrm{r}}$ rBSs are randomly chosen from the remaining 6 BSs. All selected tBSs transmit the sensing signals with equal power. Due to the absence of established data fusion algorithms for multi-target cooperative sensing within multi-antenna BS networks, we adopt the position estimation baseline using the averaging-based hard fusion approach in (85), termed Averaging.
(a) 3D Position
(b) 3D Velocity
Figure 8: Cooperative sensing accuracy versus the number of rBSs $N_{\mathrm{r}}$.
Fig. 8 shows the 3D position and velocity estimation RMSE performance of both schemes under the scenario with $K=3$ targets. A scheme termed Single-pair is also introduced as the localization performance baseline. For any target, this scheme selects the basic position estimate generated by its nearest tBS and rBS as the optimal result. Therefore, the performance of Single-pair scheme represents the upper bound of localization performance for non-cooperative bistatic sensing schemes under the considered scenario. The results in Fig. 8(a) reveal that significant accuracy improvements are achieved through multi-BS cooperation at identical transmit power levels (50 dBm or 60 dBm). Even with the minimum configuration ($N{\mathrm{t}}\times N{\mathrm{r}}=2$ tBS-rBS pairs) and the Averaging scheme, cooperative localization outperforms its non-cooperative counterpart. In addition, it can be observed from Fig. 8 that the proposed scheme substantially enhances multi-target localization accuracy compared to the baselines, thereby enabling higher 3D velocity estimation accuracy. This performance gain becomes increasingly significant with the addition of cooperative BSs, highlighting the advantages of our proposed soft fusion algorithm over traditional hard fusion methods.
(a) 3D Position
(b) 3D Velocity
Figure 9: Cooperative sensing accuracy versus transmit power.
Fig. 9 examines the impact of tBS transmit power on the RMSE of 3D position and velocity estimation. As expected, the RMSE monotonically decreases with increasing transmit power. This trend aligns with the enhanced parameter estimation accuracy observed in Fig. 6, where higher SNR improve sensing parameter extraction. Furthermore, Fig. 9 consistently demonstrates the superior positioning and velocity estimation accuracy achieved by the proposed data fusion scheme compared to the baseline. The performance gap remains significant throughout all power levels.
VI Conclusion
This paper proposed a cooperative bistatic ISAC framework for LAE applications, enabling high-accuracy multi-target sensing through standardized 5G NR infrastructure. By leveraging CP tensor decomposition with Vandermonde structural constraints, we developed a low-complexity algorithm for joint estimation of bistatic ranges, Doppler velocities, and AoAs from the multi-dimensional received echoes. To resolve data association ambiguity across distributed transceivers, an MST-based fusion scheme was designed, eliminating outliers and enabling robust 3D position and velocity reconstruction. Simulation results validated millimetre-level range accuracy and decimetre-level localization, demonstrating superior scalability for dense UAV scenarios.